OpenStudy (anonymous):

-3(1+6x) = 36-5x how do you do this with explaination

OpenStudy (anonymous):

-3(1+6x) = 36-5x First we want to get rid of the parentheses. We can do this using the distributive law -3*1 + -3*6x = 36 - 5x -3 -18x = 36 - 5x Now we want the x's on one side and the numbers (constants) on the other -18x = 36 -5x + 3 = 39 -5x -18x + 5x = 39 -13x = 39 Now we divide both sides by -13 -13x / -13 = 39 / -13 x = 39 / -13 = -3 x = -3

OpenStudy (anonymous):

thanks!